# Analyzing the dense matter equation of states in the light of the compact object HESS J1731-347

### v1

@@ Abstract
The recent mass ($0.77\pm^{0.20}_{0.17}M_{\odot}$) and radius ($10.4\pm^{0.86}_{0.78}\text{km}$) measurement of HESS J1731-347 made it one of the most fascinating object if it is indeed a neutron star. In this work, we examine the current status of the dense matter equation of states in the context of this compact object being a neutron star. We use three sets of equation of states corresponding to the three classes - neutron stars, strange stars, and hybrid stars and perform bayesian model selection on them. Our results show that for hadronic models, the EoS is preferred to be stiff at the intermediate densities. This makes the Brueckner-Hartree-Fock approximation and models based on Effective interactions deviate from current astrophysical observations on the inclusion of HESS J1731-347. Furthermore, for the strange star family, the equation of states composed of three flavor quarks prefers relatively smaller bag parameters. Analyzing the hybrid family of equation of states consisting of a first-order phase transition revealed preferences for early first-order phase transition. Comparing all the preferred equations of state among each family it was found that the current astrophysical constraints most prefer the hybrid equation of states.
@@ I Introduction
Quantum chromodynamics (QCD) predicts the existence of quarks and gluons in a deconfined state at higher densities [1, 2] hinting towards a possible phase transition (PT) at some intermediate densities [3]. The dense core of neutron stars (NSs) lies at the lower end of the intermediate densities (typical central densities lie in the range of 2-8 times that of the nuclear saturation density ($n_{s}=0.16\>\text{fm}^{-3}$), making them one of the most fascinating compact objects to study [4]. Terrestrial-based laboratories are yet to reach such densities, further fuelling our interest in these extreme objects. The core of the NSs is still a mystery, with several models suggesting the presence of only hadronic matter [5, 6, 7]. The possibility of PT at intermediate densities opens the possibility of hybrid stars [8, 9, 10, 11, 12], having an inner core composed of quark matter followed by an outer core of hadronic matter [13, 14, 15, 16, 17, 18]. There also exists another unique category of stars called strange stars (SSs) made up of strange quark matter (SQM) [19, 20, 21]. Witten suggested that the SQM is the absolute ground state consisting of $u$,$d$, and $s$ quarks supporting the idea of SSs [22]. The presence of SQM at the core of the star softens the equation of state (EoS) with possible violations in the maximum mass limit of the star imposed by PSR J0348+0432 [23] of mass $2.01\pm 0.04M_{\odot}$.
@@
However, the recent observation of HESS J1731-347 [34] started raising a few eyebrows with a mass and radius measurement of $0.77\pm^{0.20}_{0.17}M_{\odot}$ and $10.4\pm^{0.86}_{0.78}\text{km}$ respectively, making it the lightest known compact star till date. Since such low mass and radius are inherently favored by SSs, we cannot rule out the possibility of the idea of SQM inside the star. Several significant works followed this observation, each of them trying to explain the properties of this star [35, 36, 37, 38, 39]. However, one of the best way to examine any observation is to do a statistical analysis with the data (basically different models) that one has. Bayesian analysis is an important tool that is being used significantly to constrain nuclear models [40, 41, 42]. Recently, Ref [43] showed, using a bayesian framework, the impact of the compact object in the context of nuclear matter.
Bayesian inference techniques have been primarily used for constraining GW observation parameters [44, 45, 46] with Ref [47] developing a model selection technique using various astrophysical observations to compare nuclear matter EoSs. In this paper, we perform bayesian model selection to explore the implications of the compact object HESS J1731-347 along with other astrophysical observations in the light of a NS, a strange star, or a hybrid star. We consider several nuclear matter EoS models, models consisting of three flavored quarks, and also hybrid models based on first-order phase transition (FOPT). Lastly, we compare the three sets of EoSs to analyze which family of EoSs explains the current astrophysical observations the best.
The paper is arranged in the following way. Section II discusses the formalism adopted in this work to describe the EoS models and also the model selection technique and how it is implemented in this work. The results are described in section III, and finally, important results and conclusion are summarized in section IV.
@@ II.1 Hadronic EoSs
For the purpose of hadronic EoSs, we make use of the publicly available EoSs in the CompOSE repository [48]. All these EoSs consist of $npe\mu$ nuclear matter. We have considered 7 EoS models, namely the Density-dependent Relativistic Mean Field (RMF) model, the Effective Interaction model, the Thomas Fermi approach, the RMF model, the Brussels-Montreal energy density functional, non-linear RMF model, and the Brueckner-Hartree-Fock approximation. Along with these, we have also included APR(APR) (an EoS based on variational techniques), CMGO(GDFM-I) (an EoS based on density-dependent covariant density functional), and PT(GRDF2-DD2) (an EoS based on generalized relativistic density functional). More details about the EoSs used can be found in table 1.
@@
Table 1: List of all the 50 hadronic EoS with the corresponding radius of a $0.77M_{\odot}$, $1.4M_{\odot}$, and $2M_{\odot}$ star, respectively; along with the Tidal Deformability measurement of the $1.4M_{\odot}$ star and the maximum mass given by each EoS. EoSs in red are based on the density-dependent relativistic mean field (RMF) model, ones in crimson are based on Effective interactions, ones in brown are based on Thomas Fermi approach, ones in darkgreen are based on RMF approximations, ones in violet are based on Brussels-Montreal energy density functional, ones in magenta are based on non-linear RMF models and the ones in peach are based on Brueckner-Hartree-Fock approximation. The last three equations are based on variational techniques, density-dependent covariant density functional, and generalized relativistic density functional, respectively. The information of the rejected hadronic EoSs is presented with bold text.
@@ II.3 EoSs with First-Order PT
Assuming the phase transition to be of FOPT we construct the hybrid EoS from the hadronic and modified MIT bag-model quark matter EoS. The jump/transition from the hadronic phase (HP) to the quark phase (QP) happens at a particular pressure when the chemical potential of the quark phase becomes less than the chemical potential of the hadronic phase. Although this happens at a particular pressure and chemical potential, and they remain smooth throughout, there is a discontinuity in the energy density (and density) corresponding to the latent heat required for the transition. The transition is modeled using Maxwells construction as $p_{\rm HP}(\mu_{\rm c})=p_{\rm QP}(\mu_{\rm c})$ where $\mu_{\rm c}$ is the critical baryonic chemical potential where the transition occurs.
@@ II.4 Bayesian Model Selection
where $M$ refers to a model (EoS), $I$ refers to any background information that we have, and d refers to the astrophysical data. $P(M|d,I)$ is the posterior probability of the model, $P(d|M,I)$ is the marginalized-likelihood (evidence) for the data, $P(M|I)$ is the prior probability and $P(d|I)$ is a constant term.
@@
$\mathcal{O}_{M_{1}}^{M_{2}}=\dfrac{P(M_{2}|d,I)}{P(M_{1}|d,I)}=\dfrac{P(d|M_{2},I)}{P(d|M_{2},I)}\times\frac{P(M_{2}|I)}{P(M_{1}|I)}$ (3)
For uniformity, we can take the ratio of the priors to be equal to one so that $P(M_{2}|I)=P(M_{1}|I)$. By doing so, we avoid the preference for one model over the other. This choice can be modified depending on the background information.
@@
For our analysis, $d=\{d_{\rm GW},\ d_{\rm HESS},\ d_{\rm NICER}\}$ refers to the three sets of astrophysical observations we have used. The mass and tidal deformability ($\Lambda$) measurements from GW170817 [29, 30, 93] serves as $d_{\rm GW}$. For $d_{\rm NICER}$, the mass and radius measurements from PSR J0030+0451 [27] and PSR J0740+6620 [25] serve as input. Similarly, the mass and radius measurements from HESS J1731-347 [34] form $d_{\rm HESS}$. As both $d_{\rm HESS}$ and $d_{\rm NICER}$ consist only of the mass and radius measurements, the evidence calculation is the same for them.
First, let us calculate the evidence for the mass-radius measurements for the PSR J0030+0451, PSR J0740+6620, and HESS J1731-347 observations. Since we consider each EoS as a model, we replace ‘$M$’ with ‘EoS’ in eq (2). For every EoS, we solve the Tolman-Oppenheimer-Volkoff (TOV) equations [94] to obtain the mass-radius curve, also known as the MR curve. Our EoS can be parametrized either by using the mass or radius values obtained after solving the TOV equations. In our scenario, we use the mass values, and hence the evidence for the NICER observations is given by:
$P(d_{\rm NICER}|EoS,I)$ $=\>\int_{m_{min}}^{m_{max}}P(d_{\rm NICER}|m,$
@@
$\begin{split}P(m|EoS,I)&=\frac{1}{m_{max}-m_{min}};\>\text{$m_{min}\leq m\leq m_{max}$}\\ &=0\hskip 62.59596pt;\text{everywhere else}\end{split}$ (9)
@@
$\mathcal{M}_{chirp}=\dfrac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}=1.186M_{\odot}$ (11)
@@
All of the evidence integrals were performed using PyMultinest, which is a Python package for implementing the MultiNest algorithm. It offers efficient evidence calculation for multi-modal data. Furthermore, all likelihood distributions were constructed using the multivariate KDE method of Statsmodels.
@@ III Results
For our analysis, the observations we have used are $(i)$ GW170817, $(ii)$ Three X-Ray sources, namely PSR J0030+0451[27], PSR J0740+6620 [25], and HESS J1731-347 [34].
For the three sets of EoSs that we have considered in our analysis, we have 50 hadronic EoSs, 58 strange matter EoSs, and 637 hybrid EoSs. Since the hybrid EoSs were constructed using Maxwell construction, for each nuclear EoS considered, we agnostically generated a family of hybrid EoSs, resulting in a large number of hybrid EoSs. We adopt Jeffrey’s [96] scale for the log of the odds Ratio values. Note that Jeffrey’s scale is defined for the Bayes Factor and not the odds ratio, but upon taking the ratio of the prior of each model to be unity, the odds ratio becomes equal to the Bayes Factor. According to the scale, if $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-0.5, 0), then although there is evidence for model $M_{1}$, it is not worth more than a bare mention. If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-1, -0.5), then there is ‘substantial’ evidence for (against) model $M_{1}$ ($M_{2})$. If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-2, -1), then there is ‘strong’ evidence for (against) model $M_{1}$ ($M_{2}$). If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ is smaller than -2, then there is ‘decisive’ evidence for (against) model $M_{1}$ ($M_{2}$), and we can reject model $M_{2}$. Utilizing this, we present our analysis in the next subsections.
@@ III.1 Hadronic EoS
Figure 1: Odds ratio plot of SPG(M3) with other hadronic EoSs. Following Jeffrey’s scale, the region between $(-0.5,0)$ (shaded with ‘plum’) is the region in which, if a model lies, it provides evidence for the base model, but it is not worth more than a bare mention (indecisive). The region between $(-1,-0.5)$ (shaded with ‘lightcoral’) is the region of substantial evidence for the base model, and the region between $(-2,-1)$ (shaded with ‘lightpink’) is the region of strong evidence for the base model. The region beyond $-2$ (shaded with ‘mistyrose’) is the region of decisive evidence for the base model. The histogram for the odds ratio value of an EoS is depicted in the corresponding color of the respective region it lies in; for example, if an EoS is situated in the decisive region, its histogram is colored in ‘mistyrose’.
@@
After evaluating the evidence value for each EoS, the EoS with the highest evidence value is SPG(M3) (with unified NS EoS), which uses density-dependent relativistic mean-field approximation. fig. 1 shows the odds ratio plot of each EoS with respect to SPG(M3).
@@
Upon utilizing Jeffrey’s scale, as discussed before, there are exactly 23 EoS that can be decisively rejected based on their odds ratio value with respect to SPG(M3). table 1 highlights the rejected EoSs with bold text. Additionally, there are 9 EoSs with odds ratio values situated in the ‘indecisive’ region (see fig. 1 for reference). They are based on the following EoS models: density-dependent RMF model, Thomas Fermi approximation, RMF model, and generalized relativistic density functional.
@@
Figure 2: Left: EoS plot of the hadronic EoSs. The grey patch shows the region where we see a change in the stiffness of EoS. The fig. 3 shows the speed of analysis in the region of grey patch. Right: Mass-Radius curves of the hadronic EoSs. The rejected EoSs are plotted in darkblue, while the accepted ones are plotted in salmon.
@@
Using table 1 as a reference, it can be observed that except the EoSs based on Brussels-Montreal energy density functional, every other type of EoS has at least one EoS that is being rejected. However, none of the EoSs based on the Brussels-Montreal energy density functional lie in the ‘indecisive’ region. On the contrary, while one EoS based on the density-dependent RMF model can be decisively rejected, 5 (out of 7) of the EoSs lie in the ‘indecisive’ region, with SPG(M3) as the best performing EoS. This analysis suggests that the density-dependent RMF model best satisfies all the current astrophysical observations. Furthermore, it should be noted that all EoSs based on the Brueckner-Hartree-Fock approximation were rejected, leading us to the conclusion that it is the least plausible EoS type that could explain the current observations, including HESS J1731-347.
It was found that 12 (out of 16) Effective interactions based EoSs could be decisively rejected. Among the remaining four EoSs, none are situated in the comparable region; hence, among the EoS types that could not be decisively rejected, the Effective interaction type EoS is the least effective model in our analysis. Furthermore, APR(APR) (based on variational techniques) lies in the ‘strong’ region, CMGO(GDFM-1) (based on density-dependent covariant density functional) lies in the ‘strong’ region, and PT(GRDF2-DD2) (based on generalized relativistic density functional) lies in the ‘substantial’ region. In fig. 2, we show the comparison of the EoSs that were rejected along with their mass-radius curves. We see that at intermediate density ranges, the EoSs which are softer (having less speed of sound) are rejected compared to the stiffer EoSs, which is also shown in fig. 3.
@@
The adiabatic speed of sound ($c_{s}=\sqrt{\partial p/\partial\epsilon}$) is an important quantity as it determines the slope of the EoS. A stiffer EoS would tend to explain higher-mass stars. fig. 3 shows an increasing behaviour in the speed of sound which is characteristic of hadronic matter EoSs [97, 98, 99, 100, 101]. Contrary to the scenario when we do not consider the observation of HESS J1731-347 none of the hadronic EoSs used in this work could be decisively rejected as shown in fig. 4. The rejection of the EoSs from our analysis shows that all the current astrophysical observations prefer a stiffer hadronic EoS at intermediate densities.
Figure 3: The speed of sound corresponding to the shaded region in fig. 2 (left) is shown here. The y-axis represents the speed of sound in geometrized units, and the x-axis represents the energy density. The color nomenclature remains the same as previous plots.
Figure 4: Mass-Radius curves of the hadronic matter EoSs. The figure illustrates the rejected and accepted hadronic EoSs when HESS J1731-347 is not considered. Since none of the EoSs are rejected, all EoSs are plotted in salmon.

### v2

@@ Abstract
The recent mass ($0.77\pm^{0.20}_{0.17}M_{\odot}$) and radius ($10.4\pm^{0.86}_{0.78}\text{km}$) measurement of HESS J1731-347 made it one of the most fascinating object if it is indeed a neutron star. In this work, we examine the current status of the dense matter equation of states in the context of this compact object being a neutron star. We use three sets of equation of states corresponding to the three classes - neutron stars, strange stars, and hybrid stars and perform Bayesian model selection on them. Our results show that for hadronic models, the EoS is preferred to be stiff at the intermediate densities. This makes the Brueckner-Hartree-Fock approximation and models based on effective interactions deviate from current astrophysical observations on the inclusion of HESS J1731-347. Furthermore, for the strange star family, the equation of states composed of three flavor quarks prefers relatively smaller bag parameters. Analyzing the hybrid family of equation of states consisting of a first-order phase transition revealed preferences for early first-order phase transition. Comparing all the preferred equations of state among each family, it was found that the current astrophysical constraints prefer the hybrid equation of states the most.
@@ I Introduction
Quantum chromodynamics (QCD) predicts the existence of quarks and gluons in a deconfined state at higher densities [1, 2] hinting towards a possible phase transition (PT) at some intermediate densities [3]. The dense core of neutron stars (NSs) lies at the lower end of the intermediate densities (typical central densities lie in the range of 2-8 times that of the nuclear saturation density ($n_{s}=0.16\>\text{fm}^{-3}$), making them one of the most fascinating compact objects to study [4]. Terrestrial-based laboratories are yet to reach such densities, further fuelling our interest in these extreme objects.
The core of the NSs is still a mystery, with several models suggesting the presence of only hadronic matter [5, 6, 7]. The possibility of PT at intermediate densities opens the possibility of hybrid stars [8, 9, 10, 11, 12, 13], having an inner core composed of quark matter followed by an outer core of hadronic matter [14, 15, 16, 17, 18, 19]. There also exists another unique category of stars called strange stars (SSs) made up of strange quark matter (SQM) [20, 21, 22]. Witten suggested that the SQM is the absolute ground state consisting of $u$,$d$, and $s$ quarks supporting the idea of SSs [23].
@@
However, the recent observation of HESS J1731-347 [34] started raising a few eyebrows with a mass and radius measurement of $0.77\pm^{0.20}_{0.17}M_{\odot}$ and $10.4\pm^{0.86}_{0.78}\text{km}$ respectively, making it one of the lightest known compact stars till date. A recent work highlights that the analysis of the central compact object (CCO) in the supernova remnant (SNR) HESS J1731-347 is dependent on a number of necessary but insufficient conditions [35]. In particular, the authors found that the assumption of a uniform temperature carbon atmosphere (UTCA) adopted by Ref. [34] is not at all consistent with the longest & highest quality XMM-Newton data, further stating that a UTCA model for the CCO is less likely.
Although these points portray that the mass-radius calculation of the CCO is controversial, another recent work by Ref [36] talks about the minimal consistency checks for the observation of the remnant in HESS J1731-347 within existing models of SSs. They concluded with the remark that the CCO in the supernova remnant HESS J1731-347 passes initial consistency checks and can be utilised for theoretical and observational work. Furthermore, the recent work of Ref. [37] explores the possibility of HESS J1731-347 indeed being the lightest star observed. Their analysis found evidence that the star could indeed be a NS of mass 0.9 M formed from the collapse of a fast-rotating iron core. It is, therefore, necessary to at least check its effect on the present EoS bounds set by NS observations.
Several significant works followed this observation, each of them trying to explain the properties of this star [38, 39, 40, 41, 42, 43, 44, 45]. However, one of the best ways to examine any observation is to do a statistical analysis with the data (basically different models) one has. Bayesian analysis is an important tool that is being used significantly to constrain nuclear models [46, 47, 48]. Recently, Ref [49] showed, using a Bayesian framework, the impact of the compact object in the context of nuclear matter.
Bayesian inference techniques have been primarily used for constraining GW observation parameters [50, 51], with Ref [52] developing a model selection technique using various astrophysical observations to compare nuclear matter EoSs. Similarly, another paper presents the application of Bayesian model selection, ranking a few EoSs using only the GW data of GW170817 [53]. By using only GW data, one can only constrain the equation of state at densities corresponding to the central density of canonical neutron star of mass 1.4M. In this paper, we perform Bayesian model selection to explore the implications of the compact object HESS J1731-347, along with other astrophysical observations in the light of a neutron star, a strange star, or a hybrid star. Since SSs inherently favor low mass and radius, we cannot rule out the possibility of such stars. Additionally, by using several astrophysical observations, we can constrain the EoS by surveying a wide range of densities. Furthermore, we consider several nuclear matter EoS models, models consisting of three flavored quarks, and hybrid models based on first-order phase transition (FOPT). Lastly, we compare the three sets of EoSs to analyze which family of EoSs explains the current astrophysical observations the best.
The paper is arranged in the following way. Section II discusses the formalism adopted in this work to describe the EoS models and also the model selection technique and how it is implemented in this work. The results are described in section III, and finally, in section IV, we conclude with a summary and discussion of our results.
@@ II Formalism
In order to use the EoSs for astrophysical analysis, we first checked whether the EoSs satisfied constraints from chiral effective field theory. To do so, we adopted a construction similar to Refs. [33, 54], where, for densities in the range $0<n\leq 0.5n_{s}$, where $n_{s}$=0.16 fm-3, the Baym-Pethick-Sutherland (BPS) [55] EoS is considered. In the density range of $0.5n_{s}<n\leq 1.1n_{s}$, a series of monotropes of the form $p=K\rho^{\gamma}$ are constructed, where the value of ‘K’ is fixed by matching to the BPS EoS and $\gamma$ is varied in the range [1.77,3.23]. These different monotropes in the density range $0.5n_{s}<n\leq 1.1n_{s}$ form our CET band. We then check whether all the EoSs that we use for our analysis lie within this region. Our analysis did not use any EoS that did not comply with this condition. In Appendix .1, we discuss the changes in our study when the CET constraint is not taken into consideration.
@@ II.1 Hadronic EoSs
For the purpose of hadronic EoSs, we use the publicly available nuclear matter EoSs in the CompOSE repository [56]. We have considered 7 EoS models, namely the Density-dependent Relativistic Mean Field (RMF) model, the Effective Interaction model, the Thomas Fermi approach, the RMF model, the Brussels-Montreal energy density functional, non-linear RMF model, and the Brueckner-Hartree-Fock approximation. Along with these, we have also included APR(APR) (an EoS based on variational techniques), CMGO(GDFM-I) (an EoS based on density-dependent covariant density functional), and PT(GRDF2-DD2) (an EoS based on generalized relativistic density functional). More details about the EoSs used can be found in table 1.
Table 1: List of all the 38 hadronic EoS with the corresponding radius of a $0.77M_{\odot}$, $1.4M_{\odot}$, and $2M_{\odot}$ star, respectively; along with the Tidal Deformability measurement of the $1.4M_{\odot}$ star and the maximum mass given by each EoS. The last three equations are based on variational techniques, density-dependent covariant density functional, and generalized relativistic density functional, respectively. The information of the rejected hadronic EoSs is presented in bold text.
@@ II.3 EoSs with First-Order PT
Assuming the phase transition to be a FOPT, we construct the hybrid EoS from the hadronic and modified MIT bag-model quark matter EoS. The jump/transition from the hadronic phase (HP) to the quark phase (QP) happens at a particular pressure when the chemical potential of the quark phase becomes less than the chemical potential of the hadronic phase. Although this occurs at a specific pressure and chemical potential, and they remain smooth throughout, there is a discontinuity in the energy density (and density) corresponding to the latent heat required for the transition. First-order phase transition can be of two types depending on the surface tension between the adjoining fluids: Gibbs and Maxwell. The surface tension of the quark matter is the decisive parameter that dictates the type of phase transition construction mechanism. Although this parameter is poorly known, its theoretical estimates fall within a wide range of (5-300) MeV/fm3. In Refs. [92, 93] its value was found to lie in the range (10-50) MeV/fm3. However, in Ref. [94], authors estimated its value in the range (50-150) MeV/fm3 and in Ref. [95], an even higher value was estimated. Since we do not know the correct value of surface tension, both scenarios can be utilized to construct the phase transition.
In the present study, we have considered Maxwell’s construction mechanism, with $p_{\rm HP}(\mu_{\rm c})=p_{\rm QP}(\mu_{\rm c})$ (where $\mu_{\rm c}$ is the critical baryonic chemical potential where the transition occurs), while considering a large value of surface tension. It should be noted that the FOPT could have also been modeled using Gibbs construction. However, this construction requires information on the chemical potential component-wise. Since CompOSE does not provide this information, we utilize the overall chemical potential of the state, provided by CompOSE, to perform Maxwell construction.
@@ II.4 Bayesian Model Selection
where $M$ refers to a model (EoS), $I$ refers to any background information we have, and d refers to the astrophysical data. $P(M|d,I)$ is the posterior probability of the model, $P(d|M,I)$ is the marginalized-likelihood (evidence) for the data, $P(M|I)$ is the prior probability and $P(d|I)$ is a constant term.
@@
$\mathcal{O}_{M_{1}}^{M_{2}}=\dfrac{P(M_{2}|d,I)}{P(M_{1}|d,I)}=\dfrac{P(d|M_{2},I)}{P(d|M_{1},I)}\times\frac{P(M_{2}|I)}{P(M_{1}|I)}$ (3)
where the ratio of the likelihood for the two models is known as the Bayes factor. For uniformity, we can take the ratio of the priors to be equal to one so that $P(M_{2}|I)=P(M_{1}|I)$. By doing so, we avoid the preference for one model over the other. This choice can be modified depending on the background information.
@@
For our analysis, $d=\{d_{\rm GW},\ d_{\rm HESS},\ d_{\rm NICER}\}$ refers to the three sets of astrophysical observations we have used. The mass and tidal deformability ($\Lambda$) measurements from GW170817 [29, 30, 96] serves as $d_{\rm GW}$. For $d_{\rm NICER}$, the mass and radius measurements from PSR J0030+0451 [27], PSR J0437-4715, and PSR J0740+6620 [25] serve as input. Similarly, the mass and radius measurements from HESS J1731-347 [34] form $d_{\rm HESS}$. As both $d_{\rm HESS}$ and $d_{\rm NICER}$ consist only of the mass and radius measurements, the evidence calculation is the same for them.
First, let us calculate the evidence for the mass-radius measurements for the PSR J0030+0451, PSR J0740+6620, and HESS J1731-347 observations. Since we consider each EoS as a model, we replace ‘$M$’ with ‘EoS’ in eq (2). For every EoS, we solve the Tolman-Oppenheimer-Volkoff (TOV) equations [97] to obtain the mass-radius curve, also known as the MR curve. Our EoS can be parametrized either by using the mass or by the radius values obtained after solving the TOV equations. In our scenario, we use the mass values, and hence, the evidence for the NICER observations is given by:
$P(d_{\rm NICER}|{\rm EoS},I)$ $=\>\int_{m_{\rm min}}^{m_{\rm max}}P(d_{\rm NICER}|m,$
@@
$\begin{split}P(m|EoS,I)&=\frac{1}{m_{\rm max}-m_{\rm min}};\>\text{$m_{\rm min}\leq m\leq m_{\rm max}$}\\ &=0\hskip 62.59596pt;\text{everywhere else}\end{split}$ (9)
@@
$\mathcal{M}_{\rm chirp}=\dfrac{(m_{1}m_{2})^{3/5}}{(m_{1}+m_{2})^{1/5}}=1.186M_{\odot}$ (11)
@@
All of the evidence integrals were performed using PyMultiNest [99], which is a Python package for implementing the MultiNest algorithm. It offers efficient evidence calculation for multi-modal data. Furthermore, all likelihood distributions were constructed using the multivariate KDE method of Statsmodels [100].
@@ III Results
For our analysis, the observations we have used are $(i)$ GW170817, $(ii)$ Three X-Ray sources, namely PSR J0030+0451[27], PSR J0437-4715[101], PSR J0740+6620 [25], and HESS J1731-347 [34]. Appendix .2 outlines the results of our analysis when the data of HESS J1731-347 is not taken into consideration.
For the three sets of EoSs we have considered in our analysis, we have 38 hadronic EoSs, 58 strange matter EoSs, and 544 hybrid EoSs. Since the hybrid EoSs were constructed using Maxwell construction, for each nuclear EoS considered, we agnostically generated a family of hybrid EoSs, resulting in a large number of hybrid EoSs. We adopt Jeffrey’s [102] scale for the log of the odds ratio values. Jeffrey’s scale is defined for the Bayes factor, not the odds ratio. However, upon taking the ratio of the prior of each model to be unity, the odds ratio becomes equal to the Bayes factor. According to the scale, if $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-0.5, 0), then although there is evidence for model $M_{1}$, it is not worth more than a bare mention. If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-1, -0.5), then there is ‘substantial’ evidence for (against) model $M_{1}$ ($M_{2})$. If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ lies between (-2, -1), then there is ‘strong’ evidence for (against) model $M_{1}$ ($M_{2}$). If $\log_{10}\mathcal{O}_{M_{1}}^{M_{2}}$ is smaller than -2, then there is ‘decisive’ evidence for (against) model $M_{1}$ ($M_{2}$), and we can reject model $M_{2}$. Utilizing this, we present our analysis in the following subsections.
@@ III.1 Hadronic EoS
Figure 1: Odds ratio plot of XMLSLZ(DD-LZ1) with other hadronic EoSs. Following Jeffrey’s scale, the region between $(-0.5,0)$ (shaded with ‘plum’) is the region in which, if a model lies, it provides evidence for the base model, but it is not worth more than a bare mention (indecisive). The region between $(-1,-0.5)$ (shaded with ‘light coral’) is the region of substantial evidence for the base model, and the region between $(-2,-1)$ (shaded with ‘light pink’) is the region of strong evidence for the base model. The region beyond $-2$ (shaded with ‘misty rose’) is the region of decisive evidence for the base model. The histogram for the odds ratio value of an EoS is depicted in the corresponding colour of the respective region it lies in; for example, if an EoS is situated in the decisive region, its histogram is coloured in ‘misty rose’. $EoS_{i}$ refers to the equation of state being compared.
@@
After evaluating the evidence value for each EoS, the EoS with the highest evidence value is XMLSLZ(DD-LZ1), which uses the Thomas Fermi approach. Fig. 1 shows the odds ratio plot of each EoS with respect to XMLSLZ(DD-LZ1).
@@
Upon utilizing Jeffrey’s scale, as discussed before, there are precisely 15 EoS that can be decisively rejected based on their odds ratio value with respect to XMLSLZ(DD-LZ1). Table 1 highlights the rejected EoSs with bold text. Additionally, there are 9 other EoSs with odds ratio values situated in the ‘indecisive’ region (fig. 1). They are based on the following EoS models: density-dependent RMF model, Thomas Fermi approximation, RMF model, and generalized relativistic density functional.
@@
Figure 2: (a): EoS plot of the hadronic EoSs. The grey patch shows the region where there is a change in the stiffness of EoS; (b): Mass-Radius curves of the hadronic EoSs; (c): Illustrates the speed of sound corresponding to the shaded region in (a); (d): Illustrates the rejected and accepted hadronic EoSs when HESS J1731-347 is not considered. In all the figures, rejected EoSs are plotted in dark blue, while the accepted ones are plotted in salmon.
@@
Using table 1 as a reference, it can be observed that except the EoSs based on Brussels-Montreal energy density functional, every other type of EoS has at least one EoS that is being rejected. However, none of the EoSs based on the Brussels-Montreal energy density functional lie in the ‘indecisive’ region. On the contrary, while one of the EoS based on the density-dependent RMF model can be decisively rejected, 5 (out of 7) of the EoSs lie in the ‘indecisive’ region. Similarly, 3 (out of 4) of the EoSs following the Thomas Fermi approximation lie in the ‘indecisive’ region, with XMLSLZ(DD-LZ1) as the best performing EoS. This analysis suggests that the density-dependent RMF model and the Thomas Fermi approximation best satisfy all the current astrophysical observations. Additionally, the EoS PT(GRDF2-DD2) (based on generalized relativistic density functional) also performed exceptionally well.
Using table 1 and fig. 1, it can be seen that the effective-interaction based model, although not decisively rejected, is the least effective model when explaining current astrophysical bounds. Furthermore, it should be noted that all EoSs based on the Brueckner-Hartree-Fock approximation were rejected, leading us to the conclusion that it is the least plausible EoS type that could explain the current observations, including HESS J1731-347.
Most of the EoSs were accepted or rejected based on either the radius or the tidal deformability bound. Since the observations in consideration are rather recent, those EoSs that were old and did not maintain the radius or tidal deformability bounds were found to be rejected. Additionally, using table 1 and fig. 1 as a reference, it can be seen that the comparable EoSs have very similar radii for R1.4 and R2.0. We notice a trend of increasing evidence against EoSs that have increasingly dissimilar values for R1.4 and R2.0. A similar trend was also observed by Ref. [6].
In figs. 2(a) and 2(b), we show the comparison of the EoSs that were rejected along with their MR curves. Fig. 2(b) shows that the accepted EoSs have an MR curve that has a back-bending effect above 0.5 M. They also support a larger maximum mass. Additionally, the back-bending of the MR curve is reflected in the sudden stiffening of the EoS beyond $200$ MeVfm-3 in fig. 2(a). However, at much higher densities, the curves become softer. This shows that the best performing EoSs are more non-monotonous than the rejected ones.
This non-monotonous nature can be examined in greater detail in the speed of sound plot. The adiabatic speed of sound ($c_{s}=\sqrt{\partial p/\partial\epsilon}$) is an essential quantity as it determines the slope of the EoS [103, 104, 105, 106, 107]. Fig. 2(c) shows an interesting feature; the accepted curves are clearly more non-monotonic than the rejected curves, with a few even attaining a local maximum. Usually, this is associated with the production of certain new degrees of freedom or suppression of some existing degrees [108, 109, 110].
@@
Fig. 2(d) shows the rejected and accepted mass-radius curves of the hadronic EoSs when the data of HESS J1731-347 was not considered. Contrary to the scenario when we consider the observation of HESS J1731-347, none of the hadronic EoSs used in this work could be decisively rejected in this scenario, as shown in fig. 2(d). The rejection of the EoSs from our analysis shows that all the current astrophysical observations prefer a stiffer hadronic EoS at intermediate densities.